VIS Meeting on 2015/11/20
2015/10/30 13:00 - 16:00
Kamioka NAOJ Room + SeeVogh
Participants: Mark, Fabian, Takahashi, Aso, Akutsu, Hirata, Ohishi, Ishizaki, Fujii, Shoda, Okutomi, Raffaele
- Schedule status (Aso)
- Installation team organization change (Aso)
- A new team plan
- Team A: Takahashi, Shoda, Okutomi, Fujii
- Team B: Mark, Fabian, Ohishi, Shift worker ?
- Team B will start the preparation of the PR2 immediately
- A new team plan
- Test Hanging Status (Takahashi)
- BF load adjustment was done: optimal load = 53kg
EQ stop rods below the BF were wrong : EQ stop position was higher than the design -> lowered them by inserting spacers.
- Assembling IMRM: IMRM without OSEM was assembled.
- BF is already level
- Cabling: TM OSEMs and Picomotor are done. IM OSEMs, LVDT, picomotors and fishing rod needs to be done.
A cable clamp to be attached to the suspension rod has a wrong size collet. Fabian will bring the proper collets from Mitaka.
- Payload weight issue
- We need to test the theory of mistaking BF#1 and #2.
- BF load check needs to be done at the ATC
Send the BF#1 to Mitaka by 11/27 (Takahashi)
Hirata and Saito (ATC) will measure the optimal load by 12/11. Takahashi will coordinate with ATC people.
- We need BF#1 to be back to Kamioka by 12/21
- Digital system (Okutomi, Fujii)
- OSEM signal can be seen in the digital system
- Roughly checked the actuation of the mirror by OSEM by changing the offset voltage
- Damping loop of the TM is now working.
- Having a problem with excitation signal injection: waiting for Miyakawa-san to fix it.
- Working on the MEDM screen improvement
- Also modifying the realtime model of PR3
- Traverser actuation check
- Vacuum compatible motors cannot be tested in air.
- We have a vacuum non-compatible motor, so we can check the motor rotation.
Traverser cable has to be made (as a part of the cabling work)
- Clean work procedure (Aso)
Particle count record: Shoda collected data, but don't know how to extract data. Takahashi will help it.
- PR2 preparation (Mark, Fabian)
- Made an inventory of the area.
- Identified missing items (needs to be bought).
- Clean area will be ready on Nov. 25
- PR CAD model and assembly drawings (Hirata)
- Done this week: Small hexagon + BF modification. Wire clamp
Remaining: assembly drawing with the inner frame -> need the yaw rotation angle information (Aso)
- Cable status (Takahashi)
Made a cable list -> Will upload it to Wiki with more information
- Call for shift for cable checking
- One student available on the week of 11/30
Ask the student if he can come on the week of 12/7 (Takahashi)
- If the student is only available in the week of 11/30, Akutsu-kun can come to Kamioka on this week.
- Additional PR mirror box (Tatsumi):
- Waiting for delivery
- Production of additional PR winches (Akutsu)
- Ikenoue-san will try to deliver it by Dec 21.
- RM compensation mass
Design status (Fabian): working on it now, but not yet finished. Will finish sometime this weekend.
Order status (Takahashi): Haven't yet gotten a quotation. Will order early next week.
- RM EQ stop screw position (Hirata)
- ATC work (Hirata): Washed and arrived at Kamioka.
Out sourcing (Takahashi): Will ask for quotation with the RM compensation mass. All the plates are already sent to Mitaka.
- PR Optical Bench (Akutsu)
- Got half screws instead of fully threaded ones.
- Stud bolts are available from BUMAX
- SDC provided a quotation for the fully threaded screws too. Will compare these options.
- OSEM assembly status (Akutsu)
- Newly designed flexible circuits will be delivered at the end of the next week
- Next week (from 11/24) plan
- Test hanging
- Tue: Adjust the IM level.
- Wed: Level IMRM and put OSEMs, adjust the alignment of the flags. OSEM cablings
- Thu: OSEM alignment
- Fri: Signal check
- Sat: If not finished
- PR2 preparation work
- Next week: Fabian and Ohishi-san will do inventory checking, move things into the clean area.
- Next next week: Alignment of the frames.
- Test hanging
- Week of ELITES meeting (11/30-12/4)
- Mon-Thu: Alignment of the assembly and hanging frame for PR2
- Remote transfer function measurements (test hanging)
- Travel (Week of 11/24)
- Takahashi: Full week
- Shoda: Full week
- Mark: not coming
- Fabian: (Full week)
- Okutomi: Full week
- Fujii: Full week
- Travel (Week of 11/30)
- Takahashi: not coming
- Shoda: not coming
- Mark: Mon-Thu
- Fabian: not coming
- Ishizaki: Mon-Thu
- Okutomi: not coming
- Fujii: not coming
- CLIO BS
- In order to operate iKAGRA before the end of March 2016, SEO is considering the possibility of using CLIO BS instead of Type-B BS for iKAGRA.
- This idea will be discussed at the chief meeting 11/24
- BS related tasks have lower priority for the moment.
- BS Assembly Frame Design (Aso, Hirata)
- BS flag design (Mark, Fabian)
- BS gluing jigs (Tatsumi)
- Not enough crane height problem (Hirata, Saito)
- IP base clamp (Fabian)
- Bread board spring (Takahashi)
- BS clean bag (Aso)
- ETM installation procedure (Aso)
- Aso explained the basic concept of the ETM installation procedure
- A request from Mark is to make the assembly frame thinner so that we have more work space around.
Aso will compile the design information and send it to Mirapro by 11/25
- Watch dog system (Tanaka)
- no progress
- GAS blade inspection by Virgo (Takahashi)
Waiting for shipping information: Takahashi-san will send Ando-san a reminder
- Safety at the site (Takahashi, Ishizaki)
- Request to the ATC for the next year
- We need to tell ATC our requests for the next fiscal year.
- Possible tasks are:
- PRM and SRM recoil masses production: The deadline can be the end of the FY2016. Still need a rough schedule
- Type-A bottom filter production:
- Need to cut the blades. For this, we may want to check the optimal load of the SFs before deciding on the blade width of the BF.
- Checking the optimal load of SF at ATC is one of the big request candidate. We should tell ATC about this possibility.
Takahashi-san will send a plan to Akutsu-kun
- JPS meeting talks
- VIS overview: Shoda
- Mechanics: Fujii
- Control: Okutomi
- Astronomy society meeting
- Fabian will give an overview talk
2015/11/27 (Fri) 13:30 -